OpenStudy (anonymous):

integrate (1/(1+y^2)

OpenStudy (ikram002p):

\[\int\limits_{0}^{t}\frac{ 1 }{ 1+y^2 } dy \] = ?

OpenStudy (yttrium):

It's easy. It's equal to tan-1(y)+C